OpenStudy (anonymous):

find the center of mass of a thin plate of constant density covering the region bounded by parabola y= 5x-5x^2 with the line y= -5x. can anyone help please

OpenStudy (anonymous):

find the points of intersection first\[5x - 5x^{2} = -5x\]\[x^{2} - 2x = 0\]\[x \left( x-2 \right)=0\]\[x=0, x = 2\]

OpenStudy (anonymous):